# VerifiableFL: Verifiable Claims for Federated Learning using Exclaves

### v3

@@ 2.4. Approaches for verifiable claims
Given the ability of TEEs to protect computation at a hardware level and provide attestation proofs, an FL model provider could require all data providers to perform their model training inside TEEs and use remote attestation to demonstrate this in a provable manner. Existing approaches therefore propose the hosting of FL training jobs in TEEs.
@@
CFL (Guo et al., 2024) demonstrates how to host FL frameworks in TEEs based on SEV-SNP VMs coupled with NVIDIA Confidential GPUs; SecureFL (Kuznetsov et al., 2021) is an FL training system that integrates SGX-based TEEs with ARM TrustZone on the edge to ensure privacy of both models and data; PPFL (Mo et al., 2021) uses TEEs in FL training to prevent data providers from observing the model during training, thus preventing data reconstruction attacks. SecFL (Quoc and Fetzer, 2021) includes a policy manager to enforce attestation policies agreed between all clients; OpenFL (ope, 2025) uses SGX-based TEEs during FL execution to prevent model leakage; GradSec (Messaoud et al., 2021) tackles FL privacy threats by protecting a subset of the model layers inside TEEs; and CrowdGuard (Rieger et al., 2022) mitigates model poisoning by employing TEEs to perform client-side cross validation on local models.
@@
Using TEEs to provide verifiable claims about a trained FL model, however, gives rise to two challenges: (1) the attestation proofs offered by TEEs are based on static deployment time measurements of the TEE memory. The attestation proof includes a hash over the FL job code and the FL framework code and it may or may not include a hash over the datasets. This makes it challenging to relate such an attestation proof to any claims about the final trained FL model: for an auditor to verify any claims linked to such attestation proofs, the auditor would have to re-run the FL training job and observe if they obtained the same TEE attestation hashes and the same final trained FL model.
@@
In addition, (2) TEE attestation relies on the confidentiality of TEE memory because existing attestation implementations use attestation keys maintained in TEE memory and accessible by software. Since TEEs guarantee the confidentiality of TEE memory, this approach is sound unless the TEE confidentiality can be compromised. Here, TEEs have witnessed a long sequence of data leakage attacks, often by exploiting difficult to mitigate architectural side-channels (Li et al., 2022; Wang et al., 2023; Kocher et al., 2019; Li et al., 2021), which can be used to leak secrets from TEE memory. If such an attack is used successfully to leak the attestation key from a TEE, the attestation integrity can be compromised by an adversary, enabling them to manufacture bogus attestations that can be used as evidence for false model claims.
@@
Therefore, some prior work has observed that guarantees about confidentiality and integrity should be separated in trusted execution. Sealed-glass proofs (SGP) (Tramer et al., 2017) use SGX-based TEEs as “transparent enclaves” to provide attestation proofs of computation over secret inputs, while allowing a host to inspect the TEE state. Transparent enclaves, however, still assume that attestation keys remain secret: with SGX enclaves, the attestation key is held within a quoting enclave (QE) that resides in TEE memory, making it vulnerable to side-channel attacks (Van Bulck et al., 2018).
@@
Instead of relying on hardware mechanisms, other approaches have also used algorithmic and cryptographic techniques to establish claims about trained models. Prior work has proposed proof of learning approaches (Jia et al., 2021; Abbaszadeh et al., 2024), which build on techniques such as proof-of-work (PoW) (Bravo-Marquez et al., 2019), verified computation (Popa, 2024), and zero-knowledge proofs (Petkus, 2019; Garg et al., 2023). However, these approaches either incur prohibitively high overheads or can be subverted (Zhang et al., 2022; Fang et al., 2023). For example, zero-knowledge proofs such as zkSNARK (Petkus, 2019) incurs orders-of-magnitude computational overhead (Garg et al., 2023), making them impractical for ML training. PoW-based mechanisms (Bravo-Marquez et al., 2019) require the majority of the participants to follow the protocol correctly, which is an assumption that does not hold in FL, because participants are mutually untrusted and may deviate from the correct training protocol.
@@ 3. Exclaves
In this section, we introduce exclaves, a new abstraction that we use to overcome the challenges described above. We first describe the exclave concept (§3.1), and then discuss how exclaves are implemented (§3.2).
@@ 3.1. Exclave concept
To make verifiable claims about a trained FL model, one needs verifiable evidence about each of the data transformations that constituted the training process. In FL, each round of local model updates performed by a data provider would be such a data transformation; the inputs would be the current model and training data and the outputs would be the updated local model.
An exclave is an integrity-only execution environment that can be used to attest individual data transformations at runtime. By performing a data transformation within an exclave, the exclave can produce a runtime attestation proof that binds the data transformation to the specific code and data of the transformation. We refer to this attestation proof as an exclave data record (EDR).
An EDR is signed attestation proof that contains (i) a representation of a data transformation performed by that exclave, and (ii) a measurement of the exclave’s code. Formally, it is an existence proof that at least one execution of the specified code resulted in the specified data transformation.
In an EDR, the data transformation is represented by two key/value data structures: one for inputs and the other for outputs. The keys are unambiguous names for each of the inputs and outputs (e.g., model_weights or training_data). The values are fixed-size cryptographic hashes of the actual inputs and outputs. The measurement of the exclave’s code is a cryptographic hash of the code pages that were loaded into the exclave before execution began. Given access to that code, a verifier can recompute this hash and compare it against the measurement in the EDR. This is the same type of code measurement performed by current TEEs.
@@
For the measurement to remain valid throughout the life-time of the EDR, the exclave must ensure that its code is executed without external interference, and that the data within the exclave can only be modified by code within the exclave. In other words, the exclave must ensure the integrity of its code and data, even against privileged software running on the same host.
For a third party to verify that an EDR was created by a genuine exclave and has not subsequently been modified, the EDR is signed by the exclave’s attestation key. Since exclaves do not rely on data confidentiality within the exclave, i.e., they are secret-free, the attestation key cannot be stored within the exclave. As we discuss in the next section, this key should be a hardware-protected root-of-trust (HRoT) (Frazelle, 2019) stored e.g., in a separate secure co-processor. Code running inside the exclave can request the creation of an EDR, at runtime, by constructing the representation of the data transformation and sending this to the secure co-processor to create a hardware-based attestation.
@@ 3.2. Exclave implementation
By reusing existing isolation and hardware-based attestation, an exclave can be implemented with minor modifications of a TEE implementation, such as AMD SEV-SNP (Sev-Snp, 2020). Exclaves can use the hardware-based memory isolation enforced by the memory management unit (MMU) in AMD SEV-SNP: on each address translation, the MMU checks the access policy stored in the reverse map table (RMP) and maintained by the secure processor, i.e., the platform security processor (PSP) (Buhren et al., 2019). The exclave uses the PSP to generate hardware-signed attestation reports at runtime. The attestation key used for signing attestation reports, the versioned chip endorsement key (VCEK) (Buhren et al., 2019), is stored and managed by the PSP only.
Exclaves must generate runtime attestation reports by passing data securely to the PSP, without relying on shared secrets. In its current implementation, AMD SEV-SNP assumes that the TEE maintains a VM platform communication key (VMPCK) to protect communication with the PSP firmware. Since exclaves must be secret-free, the exclave implementation requires a modification of AMD SEV-SNP to protect the communication with the PSP through an MMU access check instead of a shared secret: the exclave must rely on a special protected page, accessible only by the exclave and the PSP. Supporting such a page requires a minor extension to the RMP with a new page type—an assigned page with immutable properties, owned jointly by the exclave and the PSP (via their ASIDs). The page is provisioned by the PSP on exclave creation.
Note that we are unable to prototype this change in AMD SEV-SNP because AMD does not permit the PSP to load modified firmware. In §6.3, we describe our exclave implementation on unmodified AMD SEV-SNP hardware that instead emulates the modified attestation path using a virtual TPM as a stand-in for the PSP’s attestation functionality.
@@ 4. VerifiableFL Design
Next, we describe the design of the VerifiableFL system that uses exclaves to generate EDRs as valid proofs of the FL training process. We first state our threat model (§4.1) and then present the VerifiableFL design (§4.2).
4.1. Threat model
In our threat model, an adversary may be any participant, i.e., a data provider or the model provider, and they try to invalidate claims about a trained FL model without detection by an auditor. The adversary may deliberately deviate from the correct FL training protocol by changing their storage, computation, and communicated data. We assume that the adversary has privileged software access, including control of the OS or hypervisor.
A malicious data provider may try to poison or manipulate the FL model: they may change their own training dataset; alter the FL training code or interfere with its execution at runtime; or send incorrect updates to the model provider. A malicious model provider may tamper with the aggregation code and its execution; exclude model updates by certain data providers; send messages with incorrect global models to data providers; or collude with a malicious data provider.
@@ 4.2. VerifiableFL design
The goal of the VerifiableFL design is to use exclaves to produce runtime attestation proofs, in the form of EDRs, of all executed data transformations that led to the creation of a final trained FL model. The collected set of EDRs provides a trustworthy end-to-end record of the training progress across the rounds of the FL computation. Therefore, an auditor can use the EDRs to discover violations of specific claims about the trained FL model (see §5).
@@
As shown in Fig. 3, VerifiableFL executes an FL training job as a sequence of FL tasks that define the data transformations that data providers and the model provider must execute in each training round. An FL framework such as NVFlare (Roth et al., 2022) provides the task implementation, orchestrates task execution and provides APIs for data movement between tasks. VerifiableFL executes each task within an exclave, which ensures execution integrity and produces EDRs for each executed data transformation.
@@
In a typical FL job, the tasks executed by data providers may include: (i) a local training task that takes a global FL model, reads a dataset from storage, computes a local model update and returns the updated local model; and (ii) a differential privacy (DP) task that takes a local model update, adds DP noise to protect its privacy, and outputs a DP-protected local model update. A model provider may execute: (iii) an aggregation task that receives DP-protected local model update from data providers and outputs an aggregated model; and (iv) a model update task that converts the aggregated model to a global FL model. The updated global FL model is sent to the data providers ahead of the next training round.
@@
Since VerifiableFL executes tasks inside exclaves, the produced EDRs create a trustworthy record of the executed data transformations. For example, an EDR produced by the aggregation task contains hashes of (i) the local model updates received by the task in a round and (ii) the output aggregated model; an EDR produced by the model update task has hashes of (i) the aggregated model and (ii) the updated global model; and an EDR produced by the local training task contains hashes of (i) the received global model, (ii) the training dataset and (iii) updated local model.
@@
Since the local training task reads a dataset from storage, VerifiableFL must ensure that its input hash in the EDR remains valid. Therefore, a data provider must provide the training dataset as an integrity-protected disk image (kernel development community, 2024) that uses a Merkle tree (Merkle, 1987) to ensure the integrity of the blocks in the disk image. When the dataset is accessed by the local training task in the exclave, VerifiableFL checks the dataset blocks against the Merkle tree: the root hash of the Merkle tree (combined with a random salt) serves as a non-revocable data commitment to the training dataset, which is included in the EDR.
When the FL job is executed faithfully by all participants, the EDRs must show that the correct code for the corresponding task implementation has been executed. In addition, the hashes in the EDRs that denote the data inputs and outputs must match between EDRs executed by different tasks within a training round: e.g., the output hash of the updated global model in the EDR generated by the model update task executed by the model provider must match the input hashes of the global model in the EDRs generated by the local training tasks executed by the data providers. As we describe in §5, any mismatch or inconsistencies between EDR denotes a deviation from the faithful execution of the FL job, which invalidates claims about the FL model and thus must be detected during auditing.
When the auditor finds an inconsistency between the EDR, it must be able to assign blame to a particular participant for the deviation from the faithful execution. Therefore, when an exclave generates an EDR, a participant must endorse the generated EDR by signing it with their long-term private issuer key before sending it to an EDR storage service. The EDR storage service maintains all generated EDRs during the execution of the FL job and makes them available to an auditor for verification.

### v4

@@ 2.4. Approaches for verifiable claims
Since TEEs can protect computation at a hardware level and provide attestation proofs, an FL model provider could require all data providers to perform their model training inside TEEs and use remote attestation to demonstrate this in a provable manner.
@@
Existing approaches therefore propose the hosting of FL training jobs in TEEs: CFL (guo2024trustworthy) demonstrates how to host FL frameworks in TEEs based on SEV-SNP VMs coupled with NVIDIA Confidential GPUs; SecureFL (Kuznetsov2021SecureFL) is an FL training system that integrates SGX-based TEEs with ARM TrustZone on the edge to ensure privacy of both models and data; PPFL (mo_ppfl_2021) uses TEEs to prevent data providers from observing the model during training, thus preventing data reconstruction attacks; SecFL (quoc21secfl) includes a policy manager to enforce attestation policies agreed between all clients; OpenFL (openfl) uses SGX-based TEEs during FL execution to prevent model leakage; GradSec (valerio21gradsec) targets FL privacy threats by protecting a subset of the model layers inside TEEs; and CrowdGuard (rieger_close_2022) mitigates model poisoning by employing TEEs to perform client-side cross validation on local models.
@@
Using TEEs to provide verifiable claims about a trained FL model, however, gives rise to two challenges: (1) the attestation proofs offered by TEEs are based on measurements of the TEE memory at deployment time. Such static attestation proofs includes a hash over the FL job code and the FL framework code. Typically they do not include a hash over the datasets, given that the data may be unavailable at deployment time. This makes it challenging to relate such an attestation proof to any claims about the final trained FL model: for an auditor to verify claims linked to such attestation proofs, they would have to re-run the FL training job and observe if they obtain the same final trained FL model from the same TEE attestation hashes.
@@
In addition, (2) TEE attestation relies on the confidentiality of TEE memory, because existing attestation implementations use attestation keys maintained in TEE memory. Since TEEs guarantee the confidentiality of TEE memory, this approach is sound, unless the TEE confidentiality can be compromised. Here, TEEs have witnessed a long sequence of data leakage attacks, often by exploiting difficult to mitigate architectural side-channels (li2022systematic; wang2023pwrleak; spectre; li2021cipherleaks), which can be used to leak secrets from TEE memory. If such an attack is used successfully to leak the attestation key from a TEE, the integrity of attestation proofs can be compromised by an adversary. This would enable the adversary to produce bogus attestations, which they could use as evidence for false model claims.
@@
Therefore, prior work has observed that guarantees about confidentiality and integrity should be separated in trusted execution. Sealed-glass proofs (SGP) (sgp) use SGX-based TEEs as “transparent enclaves” to provide attestation proofs of computation over secret inputs, while allowing a host to inspect the TEE memory. Transparent enclaves, however, still assume that attestation keys remain secret: with SGX enclaves, the attestation key is held inside a separate quoting enclave (QE) that resides in TEE memory, which makes it vulnerable to side-channel attacks (foreshadow).
@@
Instead of relying on hardware mechanisms, other approaches have used algorithmic and cryptographic techniques to establish claims about trained models. Proof of learning approaches (jia2021pol; abbaszadeh2024zero) build on techniques such as proof-of-work (PoW) (pow), verified computation (verified_computation), and zero-knowledge proofs (zkp1; zkp2). These approaches, however, either incur prohibitively high overheads or can be subverted (zhang2022adversarialPOL; fang2023broken). For example, zero-knowledge proofs such as zkSNARK (zkp1) introduce orders-of-magnitude computational overheads (zkp2), making them impractical for FL training. PoW-based mechanisms (pow) require the majority of the participants to follow the protocol correctly. This assumption does not hold in FL training, because participants are mutually untrusted and may deviate from the correct training protocol.
2.5. Threat model and security goals
In our threat model, we assume that any participant in the FL training job (i.e., any data provider or the centralized model provider), or any set of colluding participants, could be adversarial. This also captures the case where one or more honest participants have been compromised by an adversary and thus act adversarially.
Adversary goal. The adversary aims to deviate from the predefined FL training protocol, in a way that modifies or influences the final FL model and for this deviation to remain undetectable by other participants and/or third-party auditors.
This broad adversary goal encompasses many specific attacks that different participants could attempt: e.g., an adversarial data provider may perform fewer rounds of training to reduce resource costs; or skip a pre-processing step that would otherwise filter out specific data; or change their dataset during training to bias the final model. Similarly, an adversarial model owner may omit updates from specific data providers during the aggregation step to bias the final model; or send different models to different data providers to infer information about a specific data provider’s private dataset (boenisch2023curious). More generally, any adversarial participant may want to deviate from the predefined FL training protocol by changing their training code or by modifying or falsifying any data communicated during the FL training job.
Adversary capabilities. In an FL scenario, we assume a strong adversary with privileged access on a participant’s system. This includes full control of the OS and hypervisor of a participant, as well as the ability to intercept and modify all inbound and outbound network communication. The adversary can carry out side-channel attacks to leak the contents of a participant’s memory, even memory that is protected by a TEE. We assume, however, that the adversary will not carry out advanced physical attacks against the hardware, such as physically extracting keys from an on-chip security co-processor, due to the high costs and expertise required.
Security goals. VerifiableFL aims to ensure that the aforementioned adversary is unable to deviate from the predefined FL training protocol, in a way that modifies or influences the final model, without detection. VerifiableFL thus aims to enable any participant or third-party auditor to verify claims about the FL training process, as described in §1.
Specifically, we define a verifiable claim to be any statement about the FL training process that could have been made by a (hypothetical) trusted omniscient observer with full visibility of the training process. Claims could include statements about the executed code for training and aggregation, the number of participants and rounds of training, or any other runtime property of the FL job. It is also possible to make statements about the precise data used for training, but sometimes it is desirable for the data to remain private to the data providers. In such cases, statements can be made about a representation of the data, e.g., a blinded hash to confirm that the data has not been changed during the training process. In addition, claims may require statements about what preprocessing (e.g., filtering) has been run on the data, based on the code used.
Security non-goals. Given our assumed adversary capabilities, we cannot prevent deviation from the protocol, but we require that any such deviation is detectable. Similarly, we cannot guarantee the availability of verifiable claims, as there are various ways in which our strong adversary can prevent such claims from being made (e.g., through denial-of-service attacks). We require that, if a claim is made and verified, it must be a faithful representation of the FL training job. Finally, we do not aim to detect poisoning of a training dataset before training begins. There is a large body of related work on the detection or prevention of such attacks (rieger_close_2022). Instead, VerifiableFL can support such mitigations by (i) ensuring that they have been applied correctly by the respective participants; and (ii) allowing third parties to verify that these mitigations have been in place during training.
@@ 3. Exclaves
In this section, we introduce exclaves, a new abstraction that we use to overcome the challenges described above. We first describe the exclave concept (§3.1) and then discuss how exclaves can be implemented (§3.2).
@@ 3.1. Exclave concept
Current TEE implementations use attestation to establish a trust relationship between the TEE and a verifier. For example, an attested TLS network connection (fossati-tls-attestation-08) can prove that the verifier communicates with the TEE and that all messages received via that TLS connection are sent by the TEE.
However, as explained in §2.5, an adversary who has the ability to leak data from a TEE (e.g., through a side channel) can undermine the security guarantees of current TEE implementations, including those intended to secure FL training. In the above example, this adversary could leak the session key of the attested TLS connection and then masquerade as the TEE. In the presence of such an adversary, we cannot rely on an approach that requires data within the TEE to be kept secret from the adversary.
@@
Exclaves. Therefore, we introduce an exclave, which is an integrity-only execution environment that does not contain data that must be kept secret from the adversary. Similar to TEEs, an exclave ensures that code is executed without external interference, and that the data within the exclave can only be modified by code within the exclave. Thus, the exclave offers integrity guarantees for its code and data, even against privileged software running on the same host.
In contrast to TEEs, however, exclaves do not offer confidentiality guarantees—the only piece of data that must be kept secret from the adversary is a single attestation signing key. As we explain in §3.2, this key can be held by a secure coprocessor, which is hardened against side-channel attacks by being physically separate (frazelle2019securing).
Code running within the exclave can then request the coprocessor to produce a signed attestation report, which contains (i) a measurement of the exclave’s code and data at deployment time and (ii) additional user-defined data supplied by the code inside exclave at runtime. This type of attestation report can be used to prove that a particular statement was made by the exclave.
Exclave Data Records (EDRs). The attestation report can be used to produce an unforgeable record of data transformations carried out by the exclave (i.e., any computation inside the exclave that takes a set of data inputs and produces a set of data outputs). Exclaves can thus be used to attest individual data transformations at runtime. By performing a data transformation within an exclave, the exclave can produce a runtime attestation proof that binds the data transformation to the specific code and data of the transformation. We refer to this attestation proof as an exclave data record (EDR).
Formally, an EDR is an existence proof that at least one execution of the specified code produced the specified data transformation. In terms of its implementation, an EDR is signed statement that contains (i) a measurement of the exclave’s code, which defines the computation; and (ii) a representation of the data transformation, which contains two key/value structures: one for inputs and the other for outputs. The keys are unambiguous names for each of the inputs and outputs (e.g., model_weights or training_data). The values are fixed-size cryptographic hashes of the actual inputs and outputs. The measurement of the exclave’s code is a cryptographic hash of the code pages that were loaded into the exclave at deployment time. Given access to that code, a verifier can thus recompute this hash and compare it against the measurement in the EDR. This is the same type of code measurement performed by TEEs.
@@ 3.2. Exclave implementation
An exclave can be implemented with minor modifications of a TEE implementation, such as AMD SEV-SNP (sev2020strengthening), by reusing its existing isolation and hardware-based attestation. For example, exclaves can rely on the memory isolation enforced by the memory management unit (MMU) in AMD SEV-SNP: on each address translation, the MMU checks the access policy stored in the reverse map table (RMP) and maintained by the secure processor, i.e., the platform security processor (PSP) (buhren2019insecure). In addition, the exclave uses the PSP to generate hardware-signed attestation reports at runtime. The key used for signing attestation reports, the versioned chip endorsement key (VCEK) (buhren2019insecure), is stored and managed by the PSP only.
However, exclaves must generate runtime attestation reports by passing data securely to the PSP, without relying on shared software secrets, as this would violate their secret-free nature. Currently, AMD SEV-SNP assumes that a TEE maintains a VM platform communication key (VMPCK) to protect communication with the PSP. An exclave implementation therefore requires a modification of AMD SEV-SNP to protect the communication with the PSP through an MMU access check instead of a shared secret: the exclave must rely on a special protected page, accessible only by the exclave and the PSP, to generate attestation report securely.
Implementing exclaves using AMD-SEV SNP thus would require the following firmware changes: since an exclave would be a guest CVM running on the host, the proposed protected page could be implemented as a guest-writable analogue of SNP’s existing context page, which is RMP-protected, PSP-writable, but only guest-readable. This would require: (1) defining a new “protected” page state whose RMP entry has the same metadata as the context page, but with ownership metadata set to the exclave’s ASID; and (2) extending the access control policy to allow writes from the exclave to the protected page when the ASID matches; and (3) adding a new SNP_PROTECTED_CREATE command to provision this protected page on exclave creation (similar to context page provisioning). With these changes, access to the protected page would be enforced by SNP’s RMP check, thus enabling secret-free attestation. These firmware changes to the SNP would be feasible, because they could reuse much of the existing implementation of the context page protection.
For other TEE technologies, achieving “secret-free” attestation would require the removal of privileged software components, such as a security monitor (SM), from the runtime attestation process. For example, the Keystone TEE (keystone) would need integration with a secure hardware signer (e.g., HSM/TPM (tpm)), and its SM would become an untrusted proxy for attestation requests to that hardware signer. Compared to other TEEs, AMD SEV-SNP already provides hardware-signed runtime proofs by the PSP, which matches an exclave’s requirements.
Since we are unable to modify the proprietary firmware in AMD SEV-SNP to implement the above change to key sharing, we emulate equivalent signing functionality for exclave attestation. As we describe in §6.3, our exclave implementation on unmodified AMD SEV-SNP hardware emulates the modified attestation approach using a virtual TPM as a stand-in for the PSP’s attestation functionality with comparable performance.
@@ 4. VerifiableFL Design
Next, we describe the design of the Verifiable Federated Learning (VerifiableFL), a system that produces verifiable claims about the training process and the resulting FL model. The goal of the VerifiableFL design is to use the abstraction of exclaves to produce runtime attestation proofs, in the form of EDRs, of all executed data transformations that led to the creation of a final trained FL model. The collected set of EDRs provides a trustworthy end-to-end record of the training progress across the rounds of the FL computation. Therefore, an auditor can use the EDRs to discover violations of specific claims about the trained FL model (see §5).
The use of exclaves in the design of an FL training system introduces two challenges: (i) the integrity of the produced claims must not rely on the confidentiality of the FL training process because, unlike TEEs, exclave execution is visible to individual FL participants; and (ii) to keep the TCB small, the produced claims should only require the minimum amount of code to be executed (and thus integrity-protected) inside exclaves. In particular, the integrity of code belonging to the FL framework should not have to be protected, as long as the EDRs that attest the data transformations also establish its correct execution.
@@
As shown in Fig. 3, VerifiableFL executes an FL training job as a sequence of FL tasks that define the data transformations, which data providers and the model provider must execute in each training round. An FL framework such as NVFlare (roth2022nvidia) provides the task implementation, orchestrates task execution and provides APIs for data movement between tasks. VerifiableFL executes each task within an exclave, which ensures execution integrity and produces EDRs for each executed data transformation.
@@
In a typical FL job, the tasks executed by data providers may include: (i) a local training task that takes a global FL model, reads a dataset from storage, computes a local model update and returns the updated local model; and (ii) a differential privacy (DP) task that takes a local model update, adds DP noise to protect its privacy, and outputs a DP-protected local model update. A model provider may execute: (iii) an aggregation task that receives DP-protected local model update from data providers and outputs an aggregated model; and (iv) a model update task that converts the aggregated model to a global FL model. The updated global FL model is sent to the data providers ahead of the next training round. Note that the rest of the FL framework executes outside of exclaves, thus keeping the TCB restricted to the data transformations that impact the trained FL model.
@@
Since VerifiableFL executes tasks inside exclaves, the produced EDRs create a trustworthy record of the executed data transformations. For example, an EDR produced by the aggregation task contains hashes of (i) the local model updates received by the task in a round and (ii) the output aggregated model; an EDR produced by the model update task has hashes of (i) the aggregated model and (ii) the updated global model; and an EDR produced by the local training task contains hashes of (i) the received global model, (ii) the training dataset and (iii) updated local model. When the FL job is executed faithfully by all participants, the EDRs must show that the correct code for the task implementations has been executed. In addition, the hashes in the EDRs that denote the data inputs and outputs must match between EDRs executed by different tasks within a training round: e.g., the output hash of the updated global model in the EDR generated by the model update task executed by the model provider must match the input hashes of the global model in the EDRs generated by the local training tasks executed by the data providers.
@@
Since the local training task reads a dataset from storage, VerifiableFL must also ensure that its input hash in the EDR remains valid. Therefore, a data provider must provide the training dataset as an integrity-protected disk image (dmverity) that uses a Merkle tree (merkletree) to ensure the integrity of the blocks in the disk image. When the dataset is accessed by the local training task in the exclave, VerifiableFL checks the dataset blocks against the Merkle tree: the root hash of the Merkle tree (combined with a random salt) serves as a non-revocable data commitment to the training dataset, which is included in the EDR.
When exclaves produce EDRs, the EDRs are sent to an EDR storage service. The EDR storage service maintains all generated EDRs during the execution of the FL job and makes them available for auditing. Note that the EDR storage service itself does not need to be trusted: since EDRs are integrity-protected by a cryptographic signature, any tampering would be detected. Similarly, if individual EDRs are withheld by an adversary from the EDR storage service, the introduced inconsistencies between EDRs (see below) would reveal this.
As we describe in §5, an auditor can now check the EDRs for any mismatches or inconsistencies, which would indicate a deviation from the faithful execution of the FL job, thus invalidating claims about the trained FL model. The auditor can be an FL participant, a consumer of the trained FL model who needs to verify correct training, or a regulator who audits the training process for compliance.